the length of the portion of the normal at (1,1) to x^2=y intercepted between the axes is

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
equation of normal is given by (y-1) =( -1/2) ( x-1)
     x+ 2y -3 = 0
X intercept = 3 ; y intercept = 3/2
Length of portion intercepted = √3^2 + (3/2)^2 = 3√5/ 2.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)